OpenStudy (anonymous):

FInd the x and y interceptsfor the equation 4x-3y=9

OpenStudy (anonymous):

x-intercept is 9/4 y-intercept is -3

OpenStudy (amistre64):

still a fun, but trivial method is to divide it all by the constant, that 9 in this case. 4x - 3y = 9 ; /9 4x -3y 9 --- + --- = --- ; and simplify 9 9 9 4x -y --- + --- = 1 ; if we got a number up top, put it under 9 3 x -y ---- + --- = 1 ; if we got a "-" stick it under too 9/4 3 x y ---- + --- = 1 ; now we can just read the intercepts by 9/4 -3 the bottom numbers